# Yongtao Li, Lihua Feng, Yuejian Peng

### v1

@@ 3 Preliminaries
Furthermore, we know from Lemma 4.2 that $|E_{-}|\leq e(S)+e(T)\leq q$. By Lemma 4.2 again, we have that $|E_{+}|\leq\lfloor{n^{2}}/{4}\rfloor-e(S,T)+q\leq 5q$. By Lemma 4.4, we know that ${x}_{u}\geq 1-{(30q)}/{n}$ for every $u\in V(G)$. Therefore, we obtain
$\lambda(Y_{n,2,q})$$\geq$ $\frac{2}{\mathbf{x}^{\top}\mathbf{x}}\sum_{\{i,j\}\in E(Y_{n,2,q})}{x}_{i}{x}_{j}$
$=$ $\frac{2}{\mathbf{x}^{\top}\mathbf{x}}\sum_{\{i,j\}\in E(G)}{x}_{i}{x}_{j}+\frac{2}{\mathbf{x}^{\top}\mathbf{x}}\sum_{\{i,j\}\in E_{+}}{x}_{i}{x}_{j}-\frac{2}{\mathbf{x}^{\top}\mathbf{x}}\sum_{\{i,j\}\in E_{-}}{x}_{i}{x}_{j}$
[…]
$=$ $\lambda(G)+\frac{2}{\mathbf{x}^{\top}\mathbf{x}}\left(|E_{+}|-\frac{60q}{n}|E_{+}|+\frac{(30q)^{2}}{n^{2}}|E_{+}|-|E_{-}|\right)$
$\geq$ $\lambda(G)+\frac{2}{\mathbf{x}^{\top}\mathbf{x}}\left(1-\frac{60q}{n}|E_{+}|+\frac{(30q)^{2}}{n^{2}}|E_{+}|\right)$
@@
where the last inequality holds for $n\geq 300q^{2}$ since $|E_{+}|\leq 5q$. Therefore, we have $\lambda(Y_{n,2,q})>\lambda(G)$, as desired. ∎
From the above proof of Lemma 4.7, we can obtain the following proposition.
@@
If $n\geq 300q^{2}$ and $G$ is a graph obtained from $T_{n,2}$ by adding $q-1$ edges into the color classes arbitrarily, then $t(G)<q\lfloor{n}/{2}\rfloor$ and $\lambda(G)<\lambda(Y_{n,2,q})$.
@@
In this section, we give the proof of Theorem 2.3. Let $G$ be an $n$-vertex graph with $\lambda(G)\geq\lambda(T_{n,2,q})$ and $G$ has at most $q\lfloor{n}/{2}\rfloor$ triangles. Among such graphs, we choose $G$ as the graph with the maximum spectral radius. Our aim is to prove that $G$ is the expected extremal graph $T_{n,2,q}$. Note that $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q})$. Then Theorem 2.2 implies that $G$ has at least $q\lfloor{n}/{2}\rfloor$ triangles. Similarly, the previous Lemma 4.1 – Lemma 4.6 still hold in this setting. By Lemma 4.7, it follows that $e(G)=\lfloor n^{2}/4\rfloor+q$. By Lemma 4.2, we have $e(S)+e(T)=q$ and $e(S,T)=\lfloor n^{2}/4\rfloor$. Consequently, we know that $G$ contains exactly $q\lfloor{n}/{2}\rfloor$ triangles. Moreover, $G$ is obtained from the bipartite Turán graph $T_{n,2}$ by embedding exactly $q$ edges into the color classes, and there is no triangle among these $q$ edges. Combining these facts, it remains to show that the $q$ edges induce a star in the vertex part of size $\lceil{n}/{2}\rceil$.
Observe that $G$ does not contain $q+1$ pairwise edge-disjoint triangles. In particular, if $n$ is even and $q\neq 3$, then $|S|=|T|={n}/{2}$ and setting $k:=q+1$ in [29, Theorem 7], we know that $\lambda(G)\leq\lambda(T_{n,2,q})$, with equality if and only if $G=T_{n,2,q}$. However, the proof [29] follows the line of that of forbidding friendship graph in [9], where the triangle removal lemma is used so that it requires that $q$ is fixed and $n$ is sufficiently large. Actually, the result seems ambiguous whenever $q$ is large (say, when $q\geq\log n$). The authors [26] recently provided an alternative new method that avoids the use of triangle removal lemma.
From the above discussion, one of the main difficulties of proving Theorem 2.3 lies in giving a unique characterization of the spectral extremal graph. In the sequel, we shall give a complete proof for every $n\geq 300q^{2}$, instead of for $n$ being large enough. The key ingredient relies on the double-eigenvector technique, which was initially used by Rowlinson [50]. As far as we know, this technique turns out to be effective for treating spectral extremal graph problems; see [58, 29, 19] and references therein. The idea in our proof has its root in [29, 19]. To begin with, we need to introduce two lemmas. The first lemma provides an operation of a connected graph that increases the adjacency spectral radius strictly.
@@
Assume that $G$ is a connected graph with $u,v\in V$ and $w_{1},\ldots,w_{s}\in N(v)\setminus N(u)$. Let $\mathbf{x}=(x_{1},\ldots,x_{n})^{\top}$ be the Perron vector with $x_{v}$ corresponding to the vertex $v$. Let $G^{\prime}=G-\{vw_{i}:1\leq i\leq s\}+\{uw_{i}:1\leq i\leq s\}$. If ${x}_{u}\geq{x}_{v}$, then $\lambda(G^{\prime})>\lambda(G)$.
From the above lemma, we can get the following corollary.
Corollary 5.2.
For any $q\geq 2$, we have $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q})$.
The second lemma gives an upper bound on the sum of squares of degrees in a graph with given size. Here, we shall provide an alternative short proof.
@@
Now, we consider the case $e(H_{1})\geq 4$. Similar to the above argument, we can move the edges incident to $v^{*}$. Hence we can obtain that $d_{H_{1}}(v^{*})\geq 3$, unless $H_{1}\cong C_{4}$. For the former case, we need to show that $H_{1}$ is a star. We denote $a=d_{H_{1}}(v^{*})\geq 3$ and $H_{1}^{\prime}=H_{1}\setminus\{v^{*}\}$. If $e(H_{1}^{\prime})=0$, then $H_{1}$ is a star, so we are done. If $b:=e(H_{1}^{\prime})\geq 1$, then we define $G^{\prime}$ as the graph obtained from $G$ by deleting all $b$ edges in $H_{1}^{\prime}$, and then adding $b$ new edges from $v^{*}$ to $w_{1},\ldots,w_{b}\in S\setminus V(H_{1})$. We can see that the edges in $G^{\prime}[S]$ form a star $K_{1,a+b}$, and $G^{\prime}$ contains at most $q\lfloor{n}/{2}\rfloor$ triangles. Thus, we have $\lambda(G^{\prime})\leq\lambda(G)$. We denote $\lambda=\lambda(G)$ and $\lambda^{\prime}=\lambda\left(G^{\prime}\right)$. Let $\mathbf{y}=(y_{1},y_{2},\ldots,y_{n})^{\top}$ be the Perron vector of $G^{\prime}$. By the symmetry, we have $y_{v}=y_{w_{1}}$ for every vertex $v\in V(K_{1,a+b})\backslash\{v^{*}\}$. Thus, we have $\lambda^{\prime}y_{v^{*}}=y_{T}+(a+b)y_{w_{1}}$ and $\lambda^{\prime}y_{w_{1}}=y_{T}+y_{v^{*}}$, where $y_{T}=\sum_{u\in T}y_{u}$. Then
@@
$\mathbf{x}^{\top}\mathbf{y}\left(\lambda^{\prime}-\lambda\right)$ $=$ $\mathbf{x}^{\top}\left(A(G^{\prime})-A(G)\right)\mathbf{y}$ (2)
@@
$\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda>0.$
@@
To obtain a contradiction, we would like to show that $\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda>0$. It is sufficient to prove that $\frac{x_{v^{*}}}{x_{w_{1}}}<\frac{2a+b-{3}/{2}}{b+3}$. Note that $\lambda x_{v^{*}}=x_{T}+\sum_{u\in N_{H_{1}}(v^{*})}x_{u}\leq x_{T}+ax_{v^{*}}$. Then $x_{v^{*}}\leq\frac{x_{T}}{\lambda-a}$. Observe that $x_{w_{1}}=\frac{x_{T}}{\lambda}$. Therefore, it yields that
@@
which leads to $\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda>0$ and $\lambda^{\prime}>\lambda$, a contradiction. Thus, $H_{1}$ is a star. ∎
In Lemma 5.4, we expect to prove that $H_{i}$ is a star, instead of a $4$-cycle. But the above argument seems invalid for $H_{1}\cong C_{4}$, so we need to make an easy comparison between these two cases. If $n$ is odd, then $|S|={(n-1)}/{2}$ and $|T|={(n+1)}/{2}$. Since $G$ contains exactly $q\lfloor{n}/{2}\rfloor$ triangles, we know that the $q$ edges must be embedded into the larger part $T$. By Lemma 5.4, we must embed a star $K_{1,q}$ or a cycle $C_{4}$ (in the case of $q=4$) into $T$. By direct computation, one can check that embedding $K_{1,4}$ to $T_{n,2}$ attains the maximum spectral radius. Indeed, we can compute that $\lambda(T_{n,2,4})$ is the largest root of
@@
If $n$ is even, then $|S|=|T|={n}/{2}$. Similarly, the embedded graphs $H_{1}$ and $H_{2}$ are stars or $4$-cycles by Lemma 5.4. The remaining task is to show that embedding only one star into $S$ or $T$ will lead to a graph with the maximum spectral radius.
@@
Let $n$ be even and $G^{\prime}$ be the graph obtained from $T_{n,2}$ by embedding a star $K_{1,q}$ into the part $S$. Let $\lambda^{\prime}=\lambda(G^{\prime})$ and $\mathbf{y}$ be the Perron vector of $G^{\prime}$. Then
@@
By direct calculation, it is not hard to verify that $\lambda(G)<\lambda(T_{n,2,q})$. Next, we finish the proof in another way by using the double-eigenvector technique. We define $G^{\prime}$ to be the graph obtained from $G$ by deleting all edges of $H_{1}$ and $H_{2}$, and adding a copy $H\cong K_{1,q}$ into $S\backslash V(H_{1})$. To obtain a contradiction, it suffices to show $\lambda(G^{\prime})>\lambda(G)$. Recall that $x_{S}=\sum_{u\in S}x_{u}$ and $x_{T}=\sum_{u\in T}x_{u}$. Without loss of generality, we may assume that $x_{T}\geq x_{S}$. Let $\lambda^{\prime}=\lambda\left(G^{\prime}\right)$ and $\mathbf{y}=\left(y_{1},y_{2},\ldots,y_{n}\right)^{\top}$ be the Perron vector of $G^{\prime}$. To show $\lambda^{\prime}>\lambda$, we shall prove
$\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda\lambda^{\prime}>0.$
@@
$\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda\lambda^{\prime}$
@@
$c_{1}-c_{2}=\frac{2q_{1}q_{2}\lambda-q^{3}-q^{2}}{\lambda(\lambda-q)}>\frac{(q-1)n-q^{3}-q^{2}}{(n/2+q)n/2}>\frac{1}{10n},$
@@
$\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda\lambda^{\prime}$
@@
$\geq$ $\left(c_{2}+\frac{1}{10n}\right)x_{S}\left(y_{S}-\frac{2q}{n^{2}(n/2-q)}\lambda^{\prime}y_{V}\right)-c_{2}x_{S}y_{S}$
$>$ $\frac{x_{S}}{10n}\cdot\frac{\lambda^{\prime}y_{V}}{n+q}-(2q-1)x_{S}\cdot\frac{2q}{n^{2}(n/2-q)}\lambda^{\prime}y_{V},$
where the last inequality holds by $y_{S}=\frac{\lambda^{\prime}y_{V}}{\lambda^{\prime}+|T|}>\frac{\lambda^{\prime}y_{V}}{n+q}$ and $c_{2}<2(q-1)+\frac{q^{2}}{n/2-q}$. Hence, we can obtain $\mathbf{x}^{\top}\mathbf{y}(\lambda^{\prime}-\lambda)\cdot\lambda\lambda^{\prime}>0$ for $n\geq 300q^{2}$. This a contradiction with $\lambda^{\prime}>\lambda$. Thus, there exists exactly one $i\in\{1,2\}$ such that $H_{i}\neq\varnothing$ and then $H_{i}\cong K_{1,q}$. ∎
@@
Assume that $G$ is a graph on $n$ vertices with $\lambda(G)\geq\lambda(T_{n,2})$ and $\tau_{3}(G)\geq s$, where $s\in\mathbb{N}$. If $s=0$, then there is nothing to show. If $s=1$, then $G\neq T_{n,2}$. Theorem 1.3 implies $G$ has $\lfloor n/2\rfloor-1$ triangles. For the case $s=2$, it was proved in [27] that $G$ contains at least $n-3$ triangles. Now, we consider the case $s\geq 3$. We may assume that $t(G)\leq s{(n-1)}/{2}$. Otherwise, we are done.
Claim 1. We have $e(G)\geq\lfloor n^{2}/4\rfloor-3s$. There is a partition $V(G)=S\cup T$ such that $e(S)+e(T)<6s$, $e(S,T)>\lfloor n^{2}/4\rfloor-9s$ and ${n}/{2}-3\sqrt{s}\leq|S|,|T|\leq{n}/{2}+3\sqrt{s}$.
@@
Since $\lambda(T_{n,2})=\sqrt{\lfloor n^{2}/4\rfloor}>{(n-1)}/{2}$, by Lemma 3.2, we have
$e(G)\geq\lambda^{2}-\frac{3t}{\lambda}>\lambda^{2}-\frac{6t}{n-1}\geq\left\lfloor\frac{n^{2}}{4}\right\rfloor-3s.$
@@
Consequently, we get $e(S,T)=e(G)-e(S)-e(T)>\lfloor{n^{2}}/{4}\rfloor-9s$. By the symmetry, we may assume that $|S|\leq|T|$. Suppose on the contrary that $|S|<{n}/{2}-3\sqrt{s}$. Then $|T|=n-|S|>{n}/{2}+3\sqrt{s}$. It follows that $e(S,T)\leq|S||T|\leq\lfloor n^{2}/4\rfloor-9s$, which is a contradiction. Thus, we have ${n}/{2}-3\sqrt{s}\leq|S|,|T|\leq{n}/{2}+3\sqrt{s}$. ∎
@@
Observe that every triangle of $G$ must contain an edge from $G[S]$ or $G[T]$. If $e(S)+e(T)\leq s-1$, then all triangles of $G$ can be covered by a vertex set of at most $s-1$ vertices, contradicting with $\tau_{3}(G)\geq s$. Now, suppose that $e(S)+e(T)\geq s+1$. Note that each edge of $G[S]$ or $G[T]$ can yield at least ${n}/{2}-3\sqrt{s}$ triangles in $K_{|S|,|T|}$. From the above discussion, we know that $e(S)+e(T)<6s$ and $|S||T|-e(S,T)\leq\lfloor n^{2}/4\rfloor-e(S,T)<9s$. Then $G[S,T]$ misses less than $9s$ edges between $S$ and $T$, and each missing edge destroys at most $6s$ triangles. So $G$ contains more than $(s+1)({n}/{2}-3\sqrt{s})-9s\cdot 6s>(s+1){n}/{2}-56.5s^{2}\geq s{n}/{2}$ triangles, where the last inequality holds for every $n\geq 113s^{2}$. This contradicts with the assumption $t(G)\leq s{(n-1)}/{2}$. Thus, we have $e(S)+e(T)=s$, as claimed. ∎
From the above claim, we obtain $e(S,T)=e(G)-s>\lfloor n^{2}/4\rfloor-4s$. Then
@@
Specifically, if $s=3$, then we have $\lceil n/2\rceil-3\leq|S|,|T|\leq\lfloor n/2\rfloor+3$. Since $e(S)+e(T)=s$ and there are less than $4s$ missing edges in $G[S,T]$, we know that $G$ contains at least $s(\lceil n/2\rceil-3)-4s^{2}\geq sn/2-5s^{2}$ triangles. If $s\geq 4$, then a similar argument shows that $G$ has at least $s({n}/{2}-2\sqrt{s})-4s^{2}\geq s{n}/{2}-5s^{2}$ triangles. ∎

### v2

@@ 3.3 A spectral stability for graphs with few triangles
Furthermore, we know from Lemma 4.2 that $|E_{-}|\leq e(S)+e(T)\leq q$. By Lemma 4.3, we have that $|E_{+}|\leq\lfloor{n^{2}}/{4}\rfloor-e(S,T)+q\leq 3q$. By Lemma 4.4, we know that ${x}_{u}>1-{(20q)}/{n}$ for every $u\in V(G)$. Therefore, we obtain
$\lambda(Y_{n,2,q})$ $\geq$ $\frac{2}{\bm{x}^{\mathrm{T}}\bm{x}}\sum_{\{i,j\}\in E(Y_{n,2,q})}{x}_{i}{x}_{j}$
$=$ $\frac{2}{\bm{x}^{\mathrm{T}}\bm{x}}\sum_{\{i,j\}\in E(G)}{x}_{i}{x}_{j}+\frac{2}{\bm{x}^{\mathrm{T}}\bm{x}}\sum_{\{i,j\}\in E_{+}}{x}_{i}{x}_{j}-\frac{2}{\bm{x}^{\mathrm{T}}\bm{x}}\sum_{\{i,j\}\in E_{-}}{x}_{i}{x}_{j}$
$>$ $\lambda(G)+\frac{2}{\bm{x}^{\mathrm{T}}\bm{x}}\left(|E_{+}|\Bigl(1-\frac{20q}{n}\Bigr)^{2}-|E_{-}|\right)$
$=$ $\lambda(G)+\frac{2}{\bm{x}^{\mathrm{T}}\bm{x}}\left(|E_{+}|-\frac{40q}{n}|E_{+}|+\frac{(20q)^{2}}{n^{2}}|E_{+}|-|E_{-}|\right)$
$\geq$ $\lambda(G)+\frac{2}{\bm{x}^{\mathrm{T}}\bm{x}}\left(1-\frac{40q}{n}|E_{+}|+\frac{(20q)^{2}}{n^{2}}|E_{+}|\right)$
@@
where the last inequality holds since $n\geq 120q^{2}$ and $|E_{+}|\leq 3q$. ∎
From the above proof of Lemma 4.7, we can immediately obtain the following proposition, which can be viewed as a supplement of Example 3.3.
@@
If $q\leq\frac{1}{11}\sqrt{n}$ and $G$ is a graph obtained from $T_{n,2}$ by adding q1q-1 edges into the color classes arbitrarily, then $t(G)<q\lfloor{n}/{2}\rfloor$ and $\lambda(G)<\lambda(Y_{n,2,q})$.
@@
In this section, we are ready to prove Theorem 2.3. Equivalently, Theorem 2.3 states that if $G$ is an $n$-vertex graph with at most $q\lfloor{n}/{2}\rfloor$ triangles, then $\lambda(G)\leq\lambda(T_{n,2,q})$, with equality if and only if $G=T_{n,2,q}$. Without loss of generality, we may assume that $G$ attains the maximum spectral radius. Our goal is to prove that $G=T_{n,2,q}$. The maximality of $G$ implies $\lambda(G)>\lambda(Y_{n,2,q})$. So Theorem 2.2 is applicable, and $G$ has exactly $q\lfloor{n}/{2}\rfloor$ triangles. Similarly, the previous Lemma 4.1 – Lemma 4.6 still hold in this setting. By Lemma 4.7, it follows that $e(G)=\lfloor n^{2}/4\rfloor+q$, and $G$ is obtained from $T_{n,2}$ by embedding exactly $q$ edges into the color classes, and there is no triangle among these $q$ edges. Combining these facts, it remains to show that these $q$ edges induce a star in the vertex part of size $\lceil{n}/{2}\rceil$.
Observe that $G$ does not contain $q+1$ pairwise edge-disjoint triangles. In particular, if $n$ is even and $q\neq 3$, then |S|=|T|=n/2|S|=|T|={n}/{2} and setting $k:=q+1$ in [31, Theorem 7], we know that $\lambda(G)\leq\lambda(T_{n,2,q})$, with equality if and only if $G=T_{n,2,q}$. However, their proof in [31] follows the line of that of forbidding friendship graph in [9], where the triangle removal lemma is used so it requires that $q$ is fixed and $n$ is sufficiently large. Actually, the result seems ambiguous whenever $q$ is large (say, when $q\geq\log n$). The authors [28] recently provided an alternative new method that avoids the use of triangle removal lemma.
From the above discussion, one of the main difficulties of proving Theorem 2.3 lies in giving a unique characterization of the spectral extremal graph. In the sequel, we shall give a complete proof for every $n\geq 121q^{2}$, instead of for sufficiently large $n$. The key ingredient relies on the double-eigenvector technique, which was initially used by Rowlinson [53]. As far as we know, this technique turns out to be effective for treating spectral extremal graph problems; see [61, 31, 19, 35, 66] and references therein. The idea in our proof has its root in [31, 19]. To begin with, we need to introduce two lemmas. The first lemma provides an operation of a connected graph that strictly increases the adjacency spectral radius.
@@
Assume that $G$ is a connected graph with $u,v\in V$ and $w_{1},\ldots,w_{s}\in N(v)\setminus N(u)$. Let $\bm{x}=(x_{1},\ldots,x_{n})^{\mathrm{T}}$ be the Perron vector with $x_{v}$ corresponding to the vertex $v$. Let $G^{\prime}=G-\{vw_{i}:1\leq i\leq s\}+\{uw_{i}:1\leq i\leq s\}$. If ${x}_{u}\geq{x}_{v}$, then $\lambda(G^{\prime})>\lambda(G)$.
The second lemma gives an upper bound on the sum of squares of degrees in a graph with given size. Here, we provide a two-line simple proof for convenience.
@@
Now, we consider the case $e(H_{1})\geq 4$. Similar to the above argument, we can move the edges incident to $v^{*}$. Hence we can obtain that $d_{H_{1}}(v^{*})\geq 3$, unless $H_{1}\cong C_{4}$. For the former case, we need to show that $H_{1}$ is a star. We denote $a=d_{H_{1}}(v^{*})\geq 3$ and $H_{1}^{\prime}=H_{1}\setminus\{v^{*}\}$. If $e(H_{1}^{\prime})=0$, then $H_{1}$ is a star, so we are done. If $b:=e(H_{1}^{\prime})\geq 1$, then we define $G^{\prime}$ as the graph obtained from $G$ by deleting all $b$ edges in $H_{1}^{\prime}$, and then adding $b$ new edges from $v^{*}$ to $w_{1},\ldots,w_{b}\in S\setminus V(H_{1})$. We can see that the edges in $G^{\prime}[S]$ form a star $K_{1,a+b}$, and $G^{\prime}$ contains at most $q\lfloor{n}/{2}\rfloor$ triangles. Thus, we have $\lambda(G^{\prime})\leq\lambda(G)$. We denote $\lambda=\lambda(G)$ and $\lambda^{\prime}=\lambda\left(G^{\prime}\right)$. Let $\bm{y}=(y_{1},y_{2},\ldots,y_{n})^{\mathrm{T}}$ be the Perron vector of $G^{\prime}$. By the symmetry, we have $y_{v}=y_{w_{1}}$ for every vertex $v\in V(K_{1,a+b})\backslash\{v^{*}\}$. Thus, we have $\lambda^{\prime}y_{v^{*}}=y_{T}+(a+b)y_{w_{1}}$ and $\lambda^{\prime}y_{w_{1}}=y_{T}+y_{v^{*}}$, where $y_{T}=\sum_{u\in T}y_{u}$. Then
@@
$\bm{x}^{\mathrm{T}}\bm{y}\left(\lambda^{\prime}-\lambda\right)$ $=$ $\bm{x}^{\mathrm{T}}\left(A(G^{\prime})-A(G)\right)\bm{y}$ (2)
@@
$\bm{x}^{\mathrm{T}}\bm{y}(\lambda^{\prime}-\lambda)\cdot\lambda>0.$
@@
To obtain a contradiction, we would like to show that $\bm{x}^{\mathrm{T}}\bm{y}(\lambda^{\prime}-\lambda)\cdot\lambda>0$. It is sufficient to prove that $\frac{x_{v^{*}}}{x_{w_{1}}}<\frac{2a+b-{3}/{2}}{b+3}$. Note that $\lambda x_{v^{*}}=x_{T}+\sum_{u\in N_{H_{1}}(v^{*})}x_{u}\leq x_{T}+ax_{v^{*}}$. Then $x_{v^{*}}\leq\frac{x_{T}}{\lambda-a}$. Observe that $x_{w_{1}}=\frac{x_{T}}{\lambda}$. Therefore, it yields that
@@
which leads to $\bm{x}^{\mathrm{T}}\bm{y}(\lambda^{\prime}-\lambda)\cdot\lambda>0$ and $\lambda^{\prime}>\lambda$, a contradiction. Thus, $H_{1}$ is a star. ∎
In Lemma 5.3, we expect to prove that $H_{i}$ is a star, instead of a $4$-cycle. So we need to make an easy comparison between these two cases.
If $n$ is odd, then $|S|=\frac{n-1}{2}$ and $|T|=\frac{n+1}{2}$. Since $G$ contains exactly $q\lfloor{n}/{2}\rfloor$ triangles, the $q$ edges must be embedded into $T$. By Lemma 5.3, we must embed a star $K_{1,q}$ or a cycle $C_{4}$ (in the case of $q=4$). By computation, embedding a copy of $K_{1,4}$ attains the maximum spectral radius. Indeed, we can compute that $\lambda(T_{n,2,4})$ is the largest root of
@@
If $n$ is even, then |S|=|T|=n/2|S|=|T|={n}/{2}. Similarly, the embedded graphs $H_{1}$ and $H_{2}$ are stars or $4$-cycles by Lemma 5.3. The remaining task is to show that embedding only one star into $S$ or $T$ will lead to a graph with the maximum spectral radius.
@@
Let $n$ be even and $G^{\prime}$ be the graph obtained from $T_{n,2}$ by embedding a star $K_{1,q}$ into the part $S$. Let $\lambda^{\prime}=\lambda(G^{\prime})$ and $\bm{y}$ be the Perron vector of $G^{\prime}$. Then
@@
By direct calculation, it is not hard to verify that $\lambda(G)<\lambda(T_{n,2,q})$. Next, we finish the proof in another way by using the double-eigenvector technique. We define $G^{\prime}$ to be the graph obtained from $G$ by deleting all edges of $H_{1}$ and $H_{2}$, and adding a copy $H\cong K_{1,q}$ into $S\backslash V(H_{1})$. To obtain a contradiction, it suffices to show $\lambda(G^{\prime})>\lambda(G)$. Recall that $x_{S}=\sum_{u\in S}x_{u}$ and $x_{T}=\sum_{u\in T}x_{u}$. Without loss of generality, we may assume that $x_{T}\geq x_{S}$. Let $\lambda^{\prime}=\lambda\left(G^{\prime}\right)$ and $\bm{y}=\left(y_{1},y_{2},\ldots,y_{n}\right)^{\mathrm{T}}$ be the Perron vector of $G^{\prime}$. To show $\lambda^{\prime}>\lambda$, we shall prove
$\bm{x}^{\mathrm{T}}\bm{y}(\lambda^{\prime}-\lambda)\cdot\lambda\lambda^{\prime}>0.$
@@
$\bm{x}^{\mathrm{T}}\bm{y}(\lambda^{\prime}-\lambda)\cdot\lambda\lambda^{\prime}$
@@
$c_{1}-c_{2}=\frac{2q_{1}q_{2}\lambda-q^{3}-q^{2}}{\lambda(\lambda-q)}>\frac{(q-1)n-q^{3}-q^{2}}{(n/2+q)n/2}>\frac{1}{5n},$
@@
$\bm{x}^{\mathrm{T}}\bm{y}(\lambda^{\prime}-\lambda)\cdot\lambda\lambda^{\prime}$
@@
$\geq$ $\left(c_{2}+\frac{1}{5n}\right)x_{S}\left(y_{S}-\frac{2q}{n^{2}(n/2-q)}\lambda^{\prime}y_{V}\right)-c_{2}x_{S}y_{S}$
$>$ $\frac{x_{S}}{5n}\cdot\frac{\lambda^{\prime}y_{V}}{n+q}-(2q-1)x_{S}\cdot\frac{2q}{n^{2}(n/2-q)}\lambda^{\prime}y_{V},$
where the last inequality holds by $y_{S}=\frac{\lambda^{\prime}y_{V}}{\lambda^{\prime}+|T|}>\frac{\lambda^{\prime}y_{V}}{n+q}$ and $c_{2}<2(q-1)+\frac{q^{2}}{n/2-q}$. Hence, we can obtain $\bm{x}^{\mathrm{T}}\bm{y}(\lambda^{\prime}-\lambda)\cdot\lambda\lambda^{\prime}>0$ for $n\geq 121q^{2}$. This a contradiction with $\lambda^{\prime}>\lambda$. Thus, there exists exactly one $i\in\{1,2\}$ such that $H_{i}\neq\varnothing$ and then $H_{i}\cong K_{1,q}$. ∎
@@
Assume that $G$ is a graph on $n$ vertices with $\lambda(G)\geq\lambda(T_{n,2})$ and $\tau_{3}(G)\geq s$, where $s\geq 1$. If $s=1$, then $G\neq T_{n,2}$. Theorem 1.3 implies $G$ has $\lfloor\frac{n}{2}\rfloor-1$ triangles. For $s=2$, the authors [29] proved that $G$ has at least n3n-3 triangles. Now, we consider the case $s\geq 3$. We may assume that $t(G)\leq\frac{1}{2}s{(n-1)}$. Otherwise, we are done.
Claim 1. We have $e(G)>\lfloor n^{2}/4\rfloor-3s$. There is a partition $V(G)=S\cup T$ such that $e(S)+e(T)<6s$, $e(S,T)>\lfloor n^{2}/4\rfloor-9s$ and $\frac{n}{2}-3\sqrt{s}\leq|S|,|T|\leq\frac{n}{2}+3\sqrt{s}$.
@@
Since $\lambda(T_{n,2})=\sqrt{\lfloor n^{2}/4\rfloor}>\frac{1}{2}(n-1)$, by Lemma 3.4, we have
$e(G)\geq\lambda^{2}-\frac{3t}{\lambda}>\lambda^{2}-\frac{6t}{n-1}\geq\Big\lfloor\frac{n^{2}}{4}\Big\rfloor-3s.$
@@
Consequently, we get
$e(S,T)=e(G)-e(S)-e(T)>\Big\lfloor\frac{n^{2}}{4}\Big\rfloor-9s.$
By the symmetry, we may assume that $|S|\leq|T|$. Suppose on the contrary that $|S|<\frac{n}{2}-3\sqrt{s}$. Then $|T|=n-|S|>\frac{n}{2}+3\sqrt{s}$. It follows that $e(S,T)\leq|S||T|\leq\lfloor n^{2}/4\rfloor-9s$, which is a contradiction. Thus, we have $\frac{n}{2}-3\sqrt{s}\leq|S|,|T|\leq\frac{n}{2}+3\sqrt{s}$. ∎
@@
Observe that every triangle of $G$ must contain an edge from $G[S]$ or $G[T]$. If $e(S)+e(T)\leq s-1$, then all triangles of $G$ can be covered by a vertex set of at most s1s-1 vertices, contradicting with $\tau_{3}(G)\geq s$. Now, suppose that $e(S)+e(T)=k$ for some integer $k\geq s+1$. Note that each edge of $G[S]$ or $G[T]$ can yield at least ${n}/{2}-3\sqrt{s}$ triangles. From Claim 1, we know that $|S||T|-e(S,T)\leq\lfloor n^{2}/4\rfloor-e(S,T)<9s$. Then $G[S,T]$ misses at most (9s1)(9s-1) edges between $S$ and $T$, and each missing edge destroys at most $k$ triangles. So we get $t(G)\geq k(\frac{n}{2}-3\sqrt{s})-(9s-1)k\geq k(\frac{n}{2}-3\sqrt{s}-9s+1)>\frac{1}{2}s(n-1)$, where the last inequality holds for every $s\geq 3$ and $n\geq 28s^{2}$. This contradicts with the assumption $t(G)\leq\frac{1}{2}s(n-1)$. Thus, we have $e(S)+e(T)=s$, as claimed. ∎
From the above claims, we obtain $e(S,T)=e(G)-s>\lfloor n^{2}/4\rfloor-4s$. Then
@@
If $s=3$, then $\lceil\frac{n}{2}\rceil-3\leq|S|,|T|\leq\lfloor\frac{n}{2}\rfloor+3$. Since there are less than $4s$ missing edges in $G[S,T]$, Claim 2 implies that $t(G)\geq s(\lceil\frac{n}{2}\rceil-3)-4s\cdot s\geq\frac{1}{2}sn-5s^{2}$. If $s\geq 4$, then a similar argument shows that $t(G)\geq s(\frac{n}{2}-2\sqrt{s})-4s^{2}\geq\frac{1}{2}sn-5s^{2}$, as needed. ∎